# PTEB: Towards Robust Text Embedding Evaluation via Stochastic Paraphrasing at Evaluation Time with LLMs

### v1

@@ Abstract
Current evaluations of sentence embedding models typically rely on static test beds such as the Massive Text Embedding Benchmark (MTEB). While invaluable, repeated tuning on a fixed suite can inflate reported performance and obscure real-world robustness. We introduce the Paraphrasing Text Embedding Benchmark (PTEB), a dynamic protocol that stochastically generates meaning-preserving paraphrases at evaluation time and aggregates results across multiple runs. Using a cost-efficient LLM-based method grounded in semantic textual similarity gold ratings, we show that LLMs generate token-diverse but semantically preserving, paraphrases. Across 7 MTEB tasks, we validate our hypothesis that the performance of sentence encoders is sensitive to changes in token space even when semantics remain fixed. We also observe that smaller models are not disproportionately affected relative to larger ones. Our results are statistically robust over multiple runs and we extended our experiments to 3 multilingual datasets covering 10 languages. More generally, we aim to propose a new evaluation paradigm in NLP that relies less on static, pre-defined benchmarks but shifts towards dynamic, stochastic evaluation leveraging eval-time compute.
@@ 1 Introduction
Text embeddings have become a cornerstone of Natural Language Processing (NLP), enabling the representation of semantic meaning in dense vector spaces that can be applied to a wide range of downstream tasks. Traditionally, embeddings are evaluated on static benchmarks like the Massive Text Embedding Benchmark (MTEB) (Muennighoff et al., 2023), which has become the de facto standard.
@@
More recently, MTEB has been extended to the Massive Multilingual Text Embedding Benchmark (MMTEB) (Enevoldsen et al., 2025), covering more than 250 languages, and RTEB [^1] https://huggingface.co/blog/rteb, a retrieval benchmark with private test sets.
@@
However, benchmarks like MTEB need to be maintained and further developed (Chung et al., 2025). A persistent challenge is that static benchmarks saturate over time. Although this saturation partly reflects genuine progress, it can also be attributed to overfitting. This arises from at least two factors (Liang et al., 2025): 1. Data contamination which refers to the inclusion of benchmark items within training data, enabling models to exploit prior exposure rather than demonstrate authentic modelling capabilities (also known as data leakage). 2. Biased overtraining which involves deliberate allocation of training resources to the domains of anticipated benchmarks, resulting in models with uneven performance that excel in benchmarked domains while underperforming elsewhere. Overall, more robust embedding benchmarks are needed (Goel et al., 2025). So far, the only solutions have been temporary through private test sets or the release of new or updated benchmarks.
@@
To address this issue, we argue that model capabilities should be evaluated across dynamically generated variations of evaluation tasks at evaluation time rather than predefined static benchmarks. We therefore introduce the Paraphrasing Text Embedding Benchmark (PTEB), which generates data variations at evaluation time, creating a "population" of semantically equivalent but textually distinct problem instances that better approximates the diversity of real-world applications. Whereas MTEB targets maximal multilingual breadth and task coverage, PTEB is a complementary, compute-bounded statistically robust protocol that stress-tests semantic invariance via multi-item stochastic eval-time paraphrasing.
@@
Generative LLMs have already been shown to be effective for data augmentation through paraphrasing. In Dai et al. (2023), for example, paraphrases were used to train a BERT model (Devlin et al., 2019). Earlier, Wahle et al. (2022) demonstrated that generative LLMs can paraphrase text in a way that is difficult for humans to identify. Frank and Afli (2024) and Thirukovalluru and Dhingra (2025) used generative models to generate variations of the textual input, e.g. by paraphrasing, which are used for data augmentation at runtime to improve the performance of embeddings.
@@
Prior research on robustness and benchmarking has mainly relied on algorithmic transformations, masked language modelling (MLM), or classical word embedding techniques. For example, nlpaug (Ma, 2019) provides a general-purpose augmentation toolkit with methods such as back-translation, synonym replacement, and embedding-based substitutions (e.g., using Word2Vec (Mikolov et al., 2013)), while TextAttack (Morris et al., 2020) systematises adversarial attacks through synonym swaps, character edits, and MLM substitutions, supporting both benchmarking and adversarial training. Building on these ideas, TextFlint (Wang et al., 2021b) introduced a multilingual robustness framework combining universal and task-specific transformations incl. adversarial attacks, all validated through human evaluation. More recently, Yang et al. (2023) published an article on benchmark contamination and evasion of contamination detection methods. They demonstrated the fragility of existing evaluations for generative models, showing that rephrased benchmark samples can bypass standard decontamination methods and artificially inflate the performance of a 13b parameter model to GPT-4 levels, thus undermining benchmark trustworthiness.
@@
Whereas Yang et al. (2023) use paraphrasing to diagnose weaknesses of static benchmarks for generative LLMs, we employ eval-time paraphrasing to construct inherently contamination-resistant evaluations for embedding models. Earlier work has therefore emphasised robustness testing, augmentation, or contamination detection, but remains fundamentally constrained by static transformation rules, MLM, or classical embeddings. In contrast, our contribution is to harness state-of-the-art generative LLMs as general paraphrasers at eval-time, yielding a stochastic, dynamically moving benchmark that continuously produces novel paraphrastic test variants. This extends robustness testing from predefined perturbations to generatively rich, stochastic evaluation runs, specifically targeting embedding models across diverse NLP tasks. More broadly, our goal is to shift the focus of the NLP community from static datasets and offline augmentation to stochastic evaluation-time benchmarks powered by modern generative models.
@@ 2 Method
Figure 2: First, MTEB STS data is used to evaluate different generative LLMs for their STS rating capability. Then, using the selected LLM judge, generative LLMs are evaluated for paraphrasing MTEB STS data. Finally, the best paraphrase model is used to paraphrase MTEB datasets at eval-time generating a dynamic benchmark for embedding models.
Our method is illustrated in Figure 2. We employ generative LLMs in two distinct roles: First, as an LLM judge to select a paraphrase model. Second, as a paraphraser generating paraphrases during evaluation. To evaluate and select the LLM judge, we use sentence pairs from the MTEB STS datasets. In the absence of human evaluations or direct validation of the generated paraphrases, our method anchors PTEB in scientific rigor by grounding it in STS gold scores. Table 1 lists the STS datasets that we use and shows that the STS datasets cover a range of different text lengths ranging from short sentences to STS22’s paragraphs with an average length of 461 words. Throughout this paper, cosine similarity is employed as the similarity metric between embeddings. Generally, we use the same metrics as MTEB, for example, Spearman’s rank correlation for STS.
@@
2. Paraphrase Model Evaluation. Next, we apply generative LLMs to paraphrase the MTEB STS datasets. The LLM judge selected in the previous step rates the semantic similarity of the paraphrases on a scale from 0 to 5. Based on semantic similarity, edit distance (ED), and runtime, we select the paraphrase model ("paraphraser"). We include the ED to avoid textually overly similar paraphrases compared to the original text (larger ED is better).
3. Embedding Model Evaluation on PTEB. Finally, we employ the selected paraphraser to rephrase MTEB datasets at eval-time generating a dynamic benchmark. This allows us to evaluate various embedding models on the paraphrases and compare their scores on the original MTEB datasets. (For non-STS tasks, we use the datasets and metrics listed in Appendix A.)
@@ 3 Experiments
In this section, we present the results of our experiments. We begin by outlining the implementation details followed by the results of the LLM judge evaluation. We then examine how the selected LLM judge rates various generative models for paraphrasing. Finally, we present the results of running PTEB across different embedding models. To analyse how well the results generalise to different tasks, we evaluate datasets across all MTEB tasks and include non-English languages.
@@ 3.1 Implementation
For the LLM judge and paraphraser selection, we assessed the generative open-weights models gemma 3-27b (Gemma Team et al., 2025), gpt-oss-20b (OpenAI et al., 2025), and qwen3-32b (Yang et al., 2025). The implementation details incl. hyperparameters can be found in Appendix B.
As encoders, we evaluated qwen3-8b (Zhang et al., 2025), mxbai-embed-large-v1 (Lee et al., 2024), e5-mistral-7b-instruct (Wang et al., 2024a, b), and embeddinggemma-300m (Vera et al., 2025) as SOTA encoders. As a baseline, we included the Sentence Transformer all-mpnet-base-v2 (Reimers and Gurevych, 2020). These models were selected to include both models that currently lead or rank high on the MTEB leaderboard and a baseline model with lower performance. To ensure consistency between original and paraphrased conditions, the evaluation pipeline was re-implemented following the MTEB specification in Muennighoff et al. (2023). [^2] Our scripts are not identical with the current MTEB-evaluation but follow the original paper, e.g. MTEB now undersamples classification data to include 8 examples per label instead of using the complete train/test splits. For hyperparameters see Appendix B; for all prompts Appendix C.
3.2 Experimental Results
3.2.1 LLM Judge Evaluation
Analysing the mean absolute error (MAE) relative to the gold rating distribution (that is, the similarity scores) shows that all models have lower errors on sentence pairs with high ($\geq 4$) similarity scores Figure 3). However, while the Qwen3 and the GPT-OSS models have lower MAEs on sentence pairs with label close to 0, this does not hold for the Gemma-3 model. Table 2 presents the results for each LLM judge across datasets. Overall, gemma-3-27b achieved the highest performance on MTEB STS. With the exception of STS22 the model outperformed GPT-OSS and Qwen3 on all datasets.
Examining performance by dataset across word count bins, Figure 4 shows that sentences with $\geq 20$ words yield the lowest performance (average score 64.49%), while sentences with <10<10 words achieve an average Spearman’s rank correlation of $\geq 80\%$. In particular, the models show lower performance on long sentence pairs in STS14 and STSB with scores of 31.59% and 54.26%, respectively.
Figure 4: LLM judge scores on MTEB STS for different word count intervals; empty cells indicate that the dataset does not contain sentences with the respective word count. (in %)
Dataset gemma-3-27b gpt-oss-20b qwen-3-32b
@@ 3.2.1 LLM Judge Evaluation
Table 2: LLM judge scores on MTEB STS; best scores bold. (in %)
@@
1.9 33.28 72.74 82.76 83.18
Table 3: LLM judge ablation on the numbers of parameters for Gemma-3 grading paraphrase semantic similarity on the STS datasets. (in %)
@@ 3.2.2 Paraphrase Model Evaluation
Using gemma-3-27b as the LLM judge, we observe that also as a paraphraser it not only achieves semantic similarity scores close to the best-performing paraphraser, gpt-oss-20b, but produces paraphrases with larger ED at a lower runtime (Table 4). Hence, overall, gemma-3-27b is the most suitable paraphraser for PTEB.
3.2.3 Embedding Model Evaluation on PTEB
As single‐run estimates can be unreliable (Reimers and Gurevych, 2018; Dror et al., 2019), we evaluate STS with $n=6$ different paraphrasing seeds. Across datasets, all models decline on PTEB vs. originals with the smallest drop for embeddinggemma-300m (Table 5). e5-mistral-7b-instruct decreases by 2.70%, while others (except Gemma) drop by up to 5.09%. The gap is smallest on STS22 and largest on STS12.
Paraphrase
Model
Semantic
Sim.
Edit
Distance
Run-
time
gemma-3-27b 4.23 0.81 0.27
gpt-oss-20b 4.41 0.55 1.24
qwen3-32b 4.36 0.63 5.35
@@ 3.2.3 Embedding Model Evaluation on PTEB
Figure 5: STSB paraphrase examples.
Dataset
all-mpnet-
base-v2
embedding
gemma-300m
mxbai-embed-
large-v1
e5-mistral-
7b-instruct
qwen3-em-
bedding-8b

### v2

@@ Abstract
Current sentence embedding evaluations typically rely on static test beds like the Massive Text Embedding Benchmark (MTEB). While invaluable, repeated tuning on a fixed suite can inflate reported scores and obscure real-world robustness. We introduce the Paraphrasing Text Embedding Benchmark (PTEB), a dynamic protocol that stochastically generates meaning-preserving paraphrases at evaluation time and aggregates results across multiple runs. Using a cost-efficient LLM-based method grounded in gold ratings and human validation, we show that LLMs generate token-diverse but semantically preserving paraphrases. Across 7 MTEB tasks, we validate our hypothesis that the performance of sentence encoders is sensitive to changes in token space even when semantics remain fixed. We also observe that smaller models are not disproportionately affected relative to larger ones. Our results are statistically robust over multiple runs spanning 20 datasets and 25 languages. More generally, we aim to propose a new evaluation paradigm in NLP that relies less on static, pre-defined benchmarks but shifts towards dynamic, stochastic evaluation leveraging eval-time compute.
@@ 1 Introduction
Text embeddings have become a cornerstone of Natural Language Processing (NLP), enabling a wide range of downstream tasks. Traditionally, embeddings are evaluated on static benchmarks like the Massive Text Embedding Benchmark (MTEB) (Muennighoff et al., 2023), a de facto standard.
@@
More recently, MTEB has been extended to the Massive Multilingual Text Embedding Benchmark (MMTEB) (Enevoldsen et al., 2025), covering more than 250 languages, and RTEB [^1] https://huggingface.co/blog/rteb, a retrieval benchmark with private test sets. However, benchmarks like MTEB need to be maintained and further developed (Chung et al., 2025).
@@
A persistent challenge is that static benchmarks saturate over time. Although this saturation partly reflects genuine progress, it can also be attributed to overfitting. This arises from at least two factors (Liang et al., 2025): 1. Data contamination which refers to the inclusion of benchmark items within training data, enabling models to exploit prior exposure rather than demonstrate authentic modelling capabilities (also known as data leakage). 2. Biased overtraining which involves deliberate allocation of training resources to the domains of anticipated benchmarks, resulting in models with uneven performance that excel in benchmarked domains while underperforming elsewhere. Overall, more robust embedding benchmarks are needed (Goel et al., 2025). So far, the only solutions have been temporary through private test sets or the release of new or updated benchmarks.
@@
To address this issue, we argue that model capabilities should be evaluated on dynamically generated variations of evaluation tasks at evaluation time rather than predefined static benchmarks. We therefore introduce the Paraphrasing Text Embedding Benchmark (PTEB), which generates paraphrases at eval time, creating a set of semantically equivalent but textually distinct problem instances that better approximates diverse real-world applications. While MTEB targets maximal language and task coverage, PTEB is a complementary, statistically robust protocol that stress-tests semantic invariance via multi-run eval-time paraphrasing.
@@
Generative LLMs have been shown to be effective paraphrasers (Wahle et al., 2022), e.g., for data augmentation (Dai et al., 2023). Frank and Afli (2025) and Thirukovalluru and Dhingra (2025) used LLMs to generate variations of the input text, e.g., by paraphrasing, for data augmentation at runtime to improve the performance of embeddings.
@@
Prior research on robustness and benchmarking has mainly relied on algorithmic transformations, masked language modelling (MLM), or classical word embedding techniques. For example, nlpaug (Ma, 2019) provides a general-purpose augmentation toolkit with methods such as backtranslation, synonym replacement, and embedding-based substitutions (e.g., using Word2Vec (Mikolov et al., 2013)), while TextAttack (Morris et al., 2020) systematises adversarial attacks through synonym replacments, character edits, and MLM substitutions, supporting both benchmarking and adversarial training. Building on these ideas, TextFlint (Wang et al., 2021b) introduced a multilingual robustness framework combining universal and task-specific transformations including adversarial attacks, all validated through human evaluation.
More recently, Yang et al. (2023) published an article on benchmark contamination and evasion of contamination detection methods. They demonstrated the fragility of existing evaluations for generative models, showing that rephrased benchmark samples can bypass standard decontamination methods and artificially inflate the performance of a 13B parameter model to GPT-4 levels, thus undermining benchmark trustworthiness. Whereas Yang et al. (2023) use paraphrasing to diagnose weaknesses of static benchmarks for generative LLMs, we employ eval-time paraphrasing to construct contamination-resistant evaluations for embedding models. Li et al. (2025) address similar concerns using semantic graph manipulation to generate hard negatives by altering semantics, though limited to English. Existing work has therefore emphasised robustness testing, augmentation, contamination detection or semantic manipulation, but remains fundamentally constrained by static transformation rules, MLM, classical embeddings, and a focus on generative models or English corpora.
Concurrent work by Goel et al. (2025) proposes SAGE to analyse robustness (among other dimensions) through rules-based transformations such as synonym replacement evaluated on 3 English datasets. Our approach complements this by testing robustness to richer LLM-generated paraphrases on 20 datasets with 25 languages across six random seeds validated by LLM-judges and humans.
@@
Our contribution is to employ generative LLMs as paraphrasers at eval-time, yielding a stochastic, dynamically moving benchmark that continuously produces novel test variants. This extends robustness testing from predefined perturbations to rich, stochastic evaluation runs, specifically targeting embedding models across diverse NLP tasks. More broadly, our goal is to shift the focus of the NLP community from static datasets and offline augmentation to stochastic evaluation-time benchmarks powered by modern generative models.
@@ 2 Method
Figure 2: MTEB STS data is used to evaluate different generative LLMs for their STS rating capability. Using the selected LLM judge, generative LLMs are evaluated for paraphrasing MTEB STS data. Finally, the best paraphrase model is used to paraphrase MTEB datasets at eval-time generating a dynamic benchmark for embedding models.
Our method is illustrated in Figure 2. We apply generative LLMs in two distinct roles: First, as an LLM judge to select a paraphrase model. Second, as a paraphraser generating paraphrases during evaluation. To evaluate and select the LLM judge, we use sentence pairs from the MTEB STS datasets. Without requiring human evaluations or direct validation of the generated paraphrases, our method anchors PTEB in scientific rigor by grounding it in STS gold scores. Table 1 lists the STS datasets that we use and shows that the STS datasets cover a range of different text lengths ranging from short sentences to STS22’s paragraphs with an average length of 461 words. Throughout this paper, cosine similarity is applied as the similarity metric between embeddings. Generally, we use the same metrics as MTEB, for example, Spearman’s rank correlation for STS (see Table 13 in Appendix A for a list of all metrics per task/dataset). All averages are reported as macro-averages; for robust effect size estimation we additionally calculate the Hodges-Lehmann (HL) estimator (Hodges and Lehmann, 1963) with 95% confidence intervals (CI) for selected analyses. [^2] The HL estimator is a robust estimator calculated as the median of all Walsh averages, i.e., the pairwise means formed from every self-pair and unique unordered pair. To estimate its CIs we use a percentile bootstrap with 1000 iterations.
@@
2. Paraphrase Model Evaluation. Next, we apply generative LLMs to paraphrase the MTEB STS datasets. The LLM judge selected in the previous step rates the semantic similarity of the paraphrases on a scale from 0 to 5. Based on semantic similarity, (normalised) edit distance (ED) [^3] All reported EDs have been normalised by the maximal sentence length: $ED(s_{1},s_{2})=ED_{raw}(s_{1},s_{2})/max(len(s_{1}),len(s_{2})))$ with $s_{1}$ and $s_{2}$ being two sentences, $len(s)$ the word count of sentence $s$, and $ED_{raw}$ the non-normalised ED., and runtime, we select the paraphrase model ("paraphraser"). We include the ED as a criterion to avoid textually overly similar paraphrases compared to the original text (larger ED is better).
3. Embedding Model Evaluation on PTEB. Finally, we apply the selected paraphraser to rephrase MTEB datasets at eval-time generating a dynamic benchmark. This allows us to evaluate various embedding models on the paraphrases and compare their scores with the original MTEB datasets. For details on all datasets incl. the metrics used see Table 13 in Appendix A. We run PTEB using $n=6$ different paraphrasing seeds since single‐run estimates can be unreliable (Reimers and Gurevych, 2017, 2018; Dror et al., 2019).
Given that steps 1 and 2 of our method rely on English datasets, we validate a subset of the paraphrases generated by the selected paraphrase model for non-English datasets using human raters. For selected languages, we randomly select 50 sentence pairs each and collect 2 human ratings per pair. We established an a priori acceptability threshold of $\geq 3.00$ for the mean human similarity ratings (scale: 0 to 5). See Appendix B for further details.
@@ 3 Experiments
In this section, we present the results of our experiments. We begin by outlining the implementation details, followed by the results of the LLM judge evaluation. We then examine how the selected LLM judge rates various generative models for paraphrasing and use the best paraphraser for PTEB to evaluate various embedding models across STS and non-STS tasks as well as non-English languages. Finally, we validate PTEB using human evaluation and a prompt sensitivity analysis.
@@ 3.1 Implementation
For the LLM judge and paraphraser selection, we assessed 3 generative open-weights models: gemma 3-27b (Gemma Team et al., 2025), gpt-oss-20b (OpenAI et al., 2025), and qwen3-32b (Yang et al., 2025). Our selection criteria were: (1) open weights to enable reproducibility and auditability; (2) strong general reasoning capabilities based on public benchmarks; (3) moderate size (20-32B parameters) to run on consumer hardware with quantisation; and (4) availability via widely-used inference frameworks (Ollama, vLLM). These constraints ensure that a wide range of researchers and practitioners can reproduce and extend PTEB. Implementation details including hyperparameters can be found in Appendix C.
Dataset gemma3-27b gpt-oss-20b qwen3-32b
@@
Table 2: Spearman’s rank correlation between LLM Judge similarity scores and gold ratings for STS datasets; best scores bold. (in %)
As encoders, we evaluated qwen3-8b (Yang et al., 2025), mxbai-embed-large-v1 (Lee et al., 2024), e5-mistral-7b-instruct (Wang et al., 2024a, b), and embeddinggemma-300m (Vera et al., 2025) as SOTA encoders. As a baseline, we included the Sentence Transformer all-mpnet-base-v2 (Reimers and Gurevych, 2020). These models were selected to include both models that currently lead or rank high on the MTEB leaderboard and a baseline model with lower performance. To ensure consistency between original and paraphrased conditions, the evaluation pipeline was re-implemented following the MTEB specification in Muennighoff et al. (2023). [^4] Our scripts are not identical with the current MTEB-evaluation but follow the original paper, e.g. MTEB now undersamples classification data to include 8 examples per label instead of using the complete train/test splits. For hyperparameters see Appendix C; for all prompts Appendix D.
3.2 Experimental Results
3.2.1 LLM Judge Evaluation
Analysing the mean absolute error (MAE) relative to the gold rating distribution (that is, the similarity scores) shows that all models have lower errors on sentence pairs with high ($\geq 4$) similarity scores (Figure 3). However, while Qwen3 and GPT-OSS also have lower MAEs on sentence pairs with labels close to 0, this does not hold for the Gemma-3 model. Table 2 presents the results for each LLM judge per dataset. Overall, gemma3-27b achieved the highest performance on MTEB STS. With the exception of STS22 the model outperformed GPT-OSS and Qwen3 on all datasets. Examining performance by dataset and various word count bins (averaged over all 3 LLM judges), Figure 4 shows that sentences with $\geq 20$ words yield the lowest performance (average score 64.49%), while sentences with <10<10 words achieve an average Spearman’s rank correlation of $\geq 80\%$. In particular, on long sentence pairs in STS14, STS12, and STSB the LLM judges score only 31.59%, 52.99% and 54.26%.
Figure 4: Spearman’s rank correlation between LLM judge similarity scores and gold ratings for different word count bins; empty cells indicate that the dataset does not contain sentences with that word count. (in %)
@@ 3.2.1 LLM Judge Evaluation
1.9% 33.28% 72.74% 82.76% 83.18%
Table 3: LLM judge ablation (parameter count) for Gemma-3 grading semantic similarity on STS datasets.
@@ 3.2.2 Paraphrase Model Evaluation
Using gemma3-27b as the LLM judge, we observe that also as a paraphraser it not only achieves semantic similarity scores close to the model with the highest similarity score, gpt-oss-20b, but produces paraphrases with larger ED at a lower runtime (Table 4). The larger ED is illustrated by the examples shown in Figure 5. Hence, overall, gemma3-27b is the most suitable paraphraser for PTEB.
To mitigate potential within-model bias arising from using gemma3-27b as both the paraphraser and the LLM judge, we validated the semantic similarity ratings using gpt-oss-20b and qwen3-32b as LLM judges (Table 5). Additionally, we added mistral-small3.2:24b as a held-out judge. All non-Gemma LLM judges confirm high similarity scores for Gemma’s paraphrases.
Paraphrase Model Semantic Similarity Edit Distance Run- time
gemma3-27b 4.26 0.81 0.27
gpt-oss-20b 4.44 0.55 1.24
qwen3-32b 4.39 0.63 5.35
@@
Figure 5: Paraphrase examples (types based on Bhagat and Hovy (2013)).
LLM Judge Average HL Estimator [CI]
gemma3-27b 4.26 4.24 [4.22, 4.30]
gpt-oss-20b 4.68 4.68 [4.65, 4.72]
qwen3-32b 4.52 4.52 [4.50, 4.54]
mistral-small3.2-24b 4.22 4.22 [4.17, 4.27]
Table 5: Semantic similarity ratings (average and Hodges-Lehmann (HL) Estimator with 95% confidence intervals) between original texts and paraphrases generated with gemma--27b using different LLMs as judges.
3.2.3 PTEB for STS
Dataset all-mpnet- base-v2 embedding gemma-300m mxbai-embed- large-v1 e5-mistral- 7b-instruct qwen3-em- bedding-8b Average